a buck converter with the following parameters: Vin= 50V, D=0.4, L=270μH, C=1μF, f=25kHz, R=20Ω


a) Draw the waveform of the output voltage, Inductor current, the voltage on the diode and MOSFET. Measure the ripple in the output voltage and Inductor current.


b) Increase the switching frequency (choose any three different values) while keeping other parameters fixed (Vin= 50V, D=0.4, L=270μH, C=1μF, R=20Ω) Draw the waveform of the output voltage and Inductor current. Measure the ripple in the output voltage and Inductor current.


c) Increase the inductor size (any three different values) while keeping other parameters fixed (Vin= 50V, D=0.4, C=1μF, f=25kHz, R=20Ω) Draw the waveform of the Inductor current. Measure the ripple in the Inductor current. Comment on the results.


d) Increase the Capacitor size (any three different values) while keeping other parameters fixed (Vin= 50V, D=0.4, L=270μH, f=25kHz, R=20Ω) Draw the waveform of the output voltage. Measure the ripple in the output voltage.
Answer #1

also 1-004 Ô. 6 84625x2704 1056 af 84 276010 bx 806 (2541032 Ave - 0144 - 44.4%. — TV Nevo) DT 60-20)0.4 - 1.744 - Farbante 6for f = 80kHz . T=1 - 200sec Above derived. D= 0.4 I-D 0.6 89270x10W 1410 GOM 10332 $ SLOM 842700 10725004106 A Ven 0.1111 a